# Algorithms for the Shortest Vector Problem in $2$-dimensional Lattices, Revisited

### v1

@@ Abstract
Efficiently solving the Shortest Vector Problem (SVP) in two-dimensional lattices holds practical significance in cryptography and computational geometry. While simpler than its high-dimensional counterpart, two-dimensional SVP motivates scalable solutions for high-dimensional lattices and benefits applications like sequence cipher cryptanalysis involving large integers. In this work, we first propose a novel definition of reduced bases and develop an efficient adaptive lattice reduction algorithm CrossEuc that strategically applies the Euclidean algorithm across dimensions. Building on this framework, we introduce HVec, a vectorized generalization of the Half-GCD (HGCD) algorithm originally defined for integers, which can efficiently halve the bit-length of two vectors and may have independent interest. By iteratively invoking HVec, our optimized algorithm HVecSBP achieves a reduced basis in $O(\log nM(n))$ time for arbitrary input bases with bit-length $n$, where $M(n)$ denotes the cost of multiplying two $n$-bit integers. Compared to existing algorithms, our design is applicable to general forms of input lattices, eliminating the cost of pre-converting input bases to Hermite Normal Form (HNF). The comprehensive experimental results demonstrate that for the input lattice bases in HNF, the optimized algorithm HVecSBP achieves at least a 13.5×13.5$\times$13.5 × efficiency improvement compared to existing methods. For general-form input lattice bases, converting them to HNF before applying HVecSBP offers only marginal advantages in extreme cases where the two basis vectors are nearly degenerate (e.g., collapsing into integer-like scalars). However, as the linear dependency between input basis vectors decreases, directly employing HVecSBP yields increasingly significant efficiency gains, outperforming hybrid approaches that rely on prior HNF conversion.
@@ I Introduction
One of the most fundamental computational problems in lattice theory, known as the Shortest Vector Problem (SVP), is to find the shortest vector in a lattice. For a given lattice $\mathcal{L}$, this involves identifying a non-zero lattice vector $\mathbf{v}$ such that $||\mathbf{v}||\leq||\mathbf{u}||$ holds for all non-zero vectors $\mathbf{u}\in\mathcal{L}$. The difficulty of solving this problem grows significantly with the dimension of the lattice. However, in the simplest case of two-dimensional lattices, solving the SVP—or even the Shortest Basis Problem (SBP), which seeks a basis $\mathbf{B}=[\mathbf{a}\ \mathbf{b}]$ satisfying $\|\mathbf{a}\|=\lambda_{1}(\mathcal{L})$ and $\|\mathbf{b}\|=\lambda_{2}(\mathcal{L})$-is relatively straightforward, making it an ideal starting point for studying lattice problems. Developing fast algorithms for solving the two-dimensional SVP/SBP is of both theoretical and practical importance. From an algorithm design perspective, solving the SVP in two-dimensional lattices provides foundational insights for addressing higher-dimensional cases. For example, Lagrange’s pioneering work [5] on two-dimensional lattice reduction laid the groundwork for the celebrated LLL algorithm by A. K. Lenstra, H. W. Lenstra, and L. Lovász [6]. The LLL algorithm extends Lagrange’s approach by introducing techniques for reducing lattice bases in higher dimensions, combining ideas of approximate orthogonality and length minimization. As such, understanding and refining two-dimensional lattice reduction algorithms can directly contribute to advancements in higher-dimensional lattice computations. On the practical side, the SVP in two-dimensional lattices has direct applications in cryptoanalysis. One critical indicator of the security of cryptographic random sequences is their $2$-adic complexity, which measures the sequence’s resistance to linear feedback shift register attacks. Computing the $2$-adic complexity involves determining the Minimal Rational Fractional Representation (MRFR) of the sequence, a problem that can be directly transformed into finding the shortest basis of a two-dimensional lattice. Specifically, given a sequence, a specialized two-dimensional lattice is constructed in Hermite Normal Form (HNF), and the shortest vector in this lattice corresponds to the MRFR. Efforts to optimize algorithms for solving the MRFR problem have shown significant progress. Arnault et al. [2] proposed a Euclidean-based algorithm in 2004, with subsequent improvements in 2008 [1], to compute the shortest basis of the lattice. However, these methods occasionally produced incorrect results in certain cases. Recently, Che et al. [3] proposed a novel and faster algorithm targeting the two-dimensional lattice shortest vector problem, significantly enhancing both the computational efficiency and reliability of solutions. Further optimization of algorithms for finding the shortest basis in two-dimensional lattices would directly accelerate the computation of $2$-adic complexity, enabling faster and more accurate assessments of cryptographic random sequence security.
@@ I-A Related works
For the SVP in two-dimensional lattices, the well-known solving algorithm is the Lagrange reduction algorithm (often referred to as the Gaussian reduction algorithm) [5, 8], which is similar to the integer Euclidean algorithm. For any two lattice basis vectors $\bm{a}=[a_{1}\ a_{2}]^{T},\bm{b}=[b_{1}\ b_{2}]^{T}$, without loss of generality, we assume $\|\bm{a}\|\geq\|\bm{b}\|$, the core reduction step of Lagrange’s reduction algorithm uses a greedy approach to seek an integer
$q=\left\lceil\frac{\langle\bm{a},\bm{b}\rangle}{\langle\bm{b},\bm{b}\rangle}\right\rfloor$
that minimizes the $\ell_{2}$-length of the lattice vector $\bm{c}=\bm{a}-q\bm{b}$ to update the lattice basis vectors with $[\bm{b}\ \bm{c}]$, if the two newly obtained $\bm{b},\bm{c}$ satisfy $||\bm{b}||,||\bm{c}||\leq||\bm{b}\pm\bm{c}||$, the algorithm terminates. Otherwise, the reduction process continues using the new vectors. This algorithm can find the shortest basis with lengths $\lambda_{1}$ and $\lambda_{2}$ in time $O(M(n)n)$, where $n=\max\{\lceil\log\|\bm{a}\|\rceil,\lceil\log\|\bm{b}\|\rceil\}$ and $M(n)$ denotes the complexity of multiplying two $n$-bit integers. Clearly, the choice of $q$ is not exactly the same as that in the Euclidean algorithm of two integers. Yap [14] further studied the fully corresponding case, i.e., the algorithm gets the updated basis $[\bm{b}\ \bm{c}]$ with $\bm{c}=\bm{a}-q\bm{b}$ by taking
@@
During the process, if the update basis remains unchanged, the algorithm terminates. The author also designed a Half-Gaussian algorithm for two integer vectors, inspired by the HGCD algorithm for two integers, which improves the time complexity of lattice basis computation to $O(M(n)\log n)$. Rote’s algorithm [12] uses the lattice basis reduction method proposed by Yap [14] to solve the SVP on a two-dimensional lattice module $m$ in $O(\log m(\log\log m)^{2})$ time. Eisenbrand [4] and Che et al. [3] subsequently investigated fast algorithms for two-dimensional lattices with a known HNF basis. Specifically, when the input lattice basis satisfies $a_{1}>b_{1}\geq 0$, $a_{2}=0$ and $b_{2}\neq 0$. it was noted that by taking
@@ I-B Motivation and Our Contribution
Existing studies reveal critical gaps in solving the SVP for two-dimensional lattices:
(1) For HNF-form bases $\mathbf{B}=\begin{pmatrix}a_{1}&b_{1}\\ 0&b_{2}\end{pmatrix}$: While prior works suggest that integer HGCD algorithms can optimize continued fraction-based methods, detailed technical analyses and implementations remain scarce. As noted in [10], “the integer HGCD algorithm is intricate, error-prone, and rarely fully detailed”, and [11] emphasizes its implementation challenges due to “numerous sub-cases and limited practical adoption.” Notably, [3] provides a concrete HGCD-based SVP solution for the special case $b_{2}=1$, but the implementation details for general $b_{2}$ is not available.
(2) For general-form bases: Current algorithms first convert inputs to HNF via (HGCD-based) extended Euclidean algorithms. However, this conversion often inflates integer sizes, especially for short initial bases, leading to longer lattice vectors. This raises two interesting questions: (i) Is HNF conversion necessary? (ii) Can direct reduction algorithms bypass this step for a faster performance?
To address these gaps, this work investigates direct lattice reduction algorithms that eliminate HNF conversion and generalize HGCD-based optimizations to arbitrary $b_{2}$, bridging theoretical potential with practical implementation. Concretely, our main contributions can be summarized as follows:
@@
    Novel Definition of a Reduced Basis. For the first time, we explicitly propose a new definition of a reduced basis that differs from the conventional Lagrange-reduced basis. This novel definition expands the theoretical framework for lattice reduction, and we rigorously prove its fundamental properties. It is excepted that this lays a solid theoretical foundation for further studies in this area.
@@
    Efficient Algorithms with Detailed Implementation and Rigorous Complexity Analysis. Leveraging the newly defined reduced basis, we design a novel algorithm named CrossEuc that bypasses HNF conversion, achieving a complexity of $O(n^{2})$. This approach eliminates the computational overhead associated with HNF transformation. Furthermore, we introduce a vectorized adaptation of the HGCD algorithm, termed HVec, which represents the first explicit extension of HGCD-like techniques to vector pairs. This innovation is of independent theoretical interest and significantly advances the efficiency of solving the SVP in two-dimensional lattices. Building on this, we present an optimized algorithm HVecSBP that reduces the complexity to $O(\log n\cdot M(n))$, where $M(n)$ denotes the cost of multiplying two $n$-bit integers.
@@
    Comprehensive Experimental Performance Analysis. We conduct extensive experiments comparing our methods against state-of-the-art algorithms. Results show that (1) For the input lattice basis in HNF, our algorithm HVecSBP is about 13×13$\times$13 × faster than previous designs for the worst case. (2) For the input lattice basis in general form, compared to the previous designs without using HGCD optimization, the proposed algorithm CrossEuc achieves a at least 4.5×4.5$\times$4.5 × efficiency improvement. Compared to the previous designs using HGCD optimization, when the input vectors are nearly degenerate (i.e., treated as integers), HGCD-HNF-HVecSBP method retain a slight advantage. As the linear dependency between vectors weakens, HVecSBP exhibits remarkably growing efficiency gains.
@@ II-A Notations and Terminologies
We use $M(n)$ to denote the time complexity of the multiplication of two $n$-bit integers. Uppercase bold letters represent matrices, while lowercase bold letters represent vectors. For any two vectors $\mathbf{a}=(a_{1},a_{2})^{T}$ and $\mathbf{b}=(b_{1},b_{2})^{T}$, we define #$\#(\mathbf{a},\mathbf{b})=\max\{\#\mathbf{a},\#\mathbf{b}\}$ and #¯$\underline{\#}(\mathbf{a},\mathbf{b})=\min\{\#\mathbf{a},\#\mathbf{b}\}$. For a vector $\bm{x}=(x_{1},x_{2},\ldots,x_{m})\in\mathbb{R}^{m}$, its Euclidean length ($\ell_{2}$ norm) is defined as $\|\bm{x}\|=\sqrt{\sum_{i=1}^{m}x_{i}^{2}}$, and its Minkowski length ($\ell_{\infty}$ norm) is defined as $\|\bm{x}\|_{\infty}=\max_{1\leq i\leq m}|x_{i}|$.
@@ III Our new reduced basis and its properties
In this section, we first introduce a new reduced basis and then present its properties. Throughout this section and the rest of the paper, unless otherwise specified, the notations $\|\mathbf{x}\|$ and $\|\mathbf{x}\|_{2}$ are used to represent the Minkowski measure ($\ell_{\infty}$ norm) of $\mathbf{x}$ and the Euclidean measure ($\ell_{2}$ norm), respectively.
@@
Let $\mathcal{L}=L(\bm{B})$ be a lattice with basis $\mathbf{B}$ defined in equation (1), and let $\lambda_{1}(\mathcal{L})$ and $\lambda_{2}(\mathcal{L})$ denote the successive minims under $\ell_{2}$ norm. If $\mathbf{B}$ is reduced, then $\mathbf{u}=\mathrm{argmin}\{\|\mathbf{a}\|_{2},\|\mathbf{b}\|_{2},\|\mathbf{a}+\mathbf{b}\|_{2},\|\mathbf{a}-\mathbf{b}\|_{2}\}$ is a shortest non-zero lattice vector under the $\ell_{2}$ norm. That is, $\lambda_{1}(\mathcal{L})=\|\mathbf{u}\|_{2}$. Further, for any vector $\mathbf{x}\in\{\mathbf{a},\mathbf{b},\mathbf{a}+\mathbf{b},\mathbf{a}-\mathbf{b}\}\setminus\{\mathbf{u}\}$, $\mathbf{v}=\mathbf{x}-q\mathbf{u}$ with $q=\left\lceil\frac{\langle\mathbf{x},\mathbf{u}\rangle}{\langle\mathbf{u},\mathbf{u}\rangle}\right\rfloor$ is a lattice vector achieving $\lambda_{2}(\mathcal{L})$. That is, $\lambda_{2}(\mathcal{L})=\|\mathbf{v}\|_{2}$.

### v2

@@ Abstract
Efficiently solving the Shortest Vector (Basis) Problem in two-dimensional lattices holds practical significance in cryptography and computational geometry. While simpler than its high-dimensional counterpart, two-dimensional SVP motivates scalable solutions for high-dimensional lattices and benefits applications like sequence cipher cryptanalysis involving large integers. In this work, we first propose a novel definition of reduced bases and develop an efficient adaptive lattice reduction algorithm CrossEuc that strategically applies the Euclidean algorithm across dimensions. Building on this framework, we introduce HVec, a vectorized generalization of the integer Half-GCD (HGCD) algorithm that efficiently halves vector bit-lengths, which may be of independent interest. By iteratively invoking HVec, our optimized algorithm HVecSBP achieves a reduced basis in $O(\log nM(n))$ time for arbitrary input bases with bit-length $n$, where $M(n)$ denotes the cost of multiplying two $n$-bit integers.
For finding the $\ell_{2}$-shortest basis, our new algorithms avoids costly inner products in its basic reduction step. The comprehensive experimental results demonstrate that (1) $\mathbf{CrossEuc}$ achieves approximately 500×500$\times$500 × and 350×350$\times$350 × speedup compared to Yap’s foundamental reduction algorithm $\mathbf{CRS}$ and the classical Lagrange reduction algorithm, respectively. (2) our optimized variant, $\mathbf{HVecSBP}$, delivers an additional 14$\times$ acceleration over Yap’s optimized $\mathbf{HalfGaussianSBP}$. For finding the $\ell_{\infty}$-shortest basis, our new designs eliminate Hermite Normal Form (HNF) conversion costs for general-form inputs and experimentally demonstrate significant improvements: (1) For HNF-form bases, $\mathbf{HVecSBP}$ achieves 13×13$\times$13 × acceleration over prior designs in worst-case scenarios; (2) For general-form bases,$\mathbf{CrossEuc}$ attains $\geq$4.5$\times$ efficiency gains over non-HGCD methods. Crucially, $\mathbf{HVecSBP}$ exhibits progressively intensifying advantages against HGCD-optimized designs as lattice basis vectors linear dependence weakens.
@@ I Introduction
One of the most fundamental computational problems in lattice theory, known as the Shortest Vector Problem (SVP), is to find the shortest vector in a lattice. For a given lattice $\mathcal{L}$, this involves identifying a non-zero lattice vector $\mathbf{v}$ such that $||\mathbf{v}||\leq||\mathbf{u}||$ holds for all non-zero vectors $\mathbf{u}\in\mathcal{L}$. The difficulty of solving this problem grows significantly with the dimension of the lattice. However, in the simplest case of two-dimensional lattices, solving the SVP—or even the Shortest Basis Problem (SBP), which seeks a basis $\mathbf{B}=[\mathbf{a}\ \mathbf{b}]$ satisfying $\|\mathbf{a}\|=\lambda_{1}(\mathcal{L})$ and $\|\mathbf{b}\|=\lambda_{2}(\mathcal{L})$-is relatively straightforward, making it an ideal starting point for studying lattice problems. Developing fast algorithms for solving the two-dimensional SVP/SBP is of both theoretical and practical importance. From an algorithm design perspective, solving the SVP in two-dimensional lattices provides foundational insights for addressing higher-dimensional cases. For example, Lagrange’s pioneering work [6] on two-dimensional lattice reduction laid the groundwork for the celebrated LLL algorithm by A. K. Lenstra, H. W. Lenstra, and L. Lovász [7]. The LLL algorithm extends Lagrange’s approach by introducing techniques for reducing lattice bases in higher dimensions, combining ideas of approximate orthogonality and length minimization. As such, understanding and refining two-dimensional lattice reduction algorithms can directly contribute to advancements in higher-dimensional lattice computations. On the practical side, the SVP in two-dimensional lattices has direct applications in cryptoanalysis. One critical indicator of the security of cryptographic random sequences is their $2$-adic complexity, which measures the sequence’s resistance to linear feedback shift register attacks. Computing the $2$-adic complexity involves determining the Minimal Rational Fractional Representation (MRFR) of the sequence, a problem that can be directly transformed into finding the shortest basis of a two-dimensional lattice in the $\ell_{\infty}$ norm. Specifically, given a sequence, a specialized two-dimensional lattice is constructed in Hermite Normal Form (HNF), and the shortest vector in this lattice corresponds to the MRFR. Efforts to optimize algorithms for solving the MRFR problem have shown significant progress. Arnault et al. [2] proposed a Euclidean-based algorithm in 2004, with subsequent improvements in 2008 [1], to compute the shortest basis of the lattice. However, these methods occasionally produced incorrect results in certain cases. Recently, Che et al. [3] proposed a novel and faster algorithm targeting the two-dimensional lattice shortest vector problem, significantly enhancing both the computational efficiency and reliability of solutions. Further optimization of algorithms for finding the shortest basis in two-dimensional lattices would directly accelerate the computation of $2$-adic complexity, enabling faster and more accurate assessments of cryptographic random sequence security.
@@ I-A Related works
[…]
to minimize the length of $\bm{c}=\bm{a}-q\bm{b}$, updating the basis to $[\bm{b},\bm{c}]$. Iterating this process yields the shortest basis in $O(M(n)n)$ time, where $n=\max\{\lceil\log\|\bm{a}\|\rceil,\lceil\log\|\bm{b}\|\rceil\}$ and $M(n)$ is the complexity of multiplying two $n$-bit integers.
For the frequently used Euclidean metric ($\ell_{2}$ norm), Lagrange reduction uses:
$q=\left\lceil\frac{\langle\bm{a},\bm{b}\rangle}{\langle\bm{b},\bm{b}\rangle}\right\rfloor.$
This $q$-selection differs from the standard integer Euclidean algorithm. Yap [15] studied a directly analogous approach, updating the basis to $[\bm{b},\bm{c}]$ with $\bm{c}=\bm{a}-q\bm{b}$ via:
@@
The algorithm terminates if the basis remains unchanged after an update. This method, termed $\mathbf{CRS}$, was proven to have slightly higher complexity than Lagrange’s reduction, but with negligible efficiency loss (”not giving up too much” as said in [15]). Crucially, it preserves the geometric property that basis vectors always form acute angles. Further optimizing $\mathbf{CRS}$, Yap designed a Half-Gaussian algorithm for integer vectors inspired by the integer HGCD method, reducing the time complexity to $O(M(n)\log n)$. Rote’s algorithm [13] leverages Yap’s lattice basis reduction [15] to solve SVP for a two-dimensional lattice module $m$ in $O(\log m(\log\log m)^{2})$ time.
For the special Minkowski metric ($\ell_{\infty}$ norm), Eisenbrand [4] and Che et al. [3] subsequently investigated fast algorithms for two-dimensional lattices with a known HNF basis. Specifically, when the input lattice basis satisfies $a_{1}>b_{1}\geq 0$, $a_{2}=0$ and $b_{2}\neq 0$. it was noted that by taking
@@
Regarding the relationship between the shortest vector under the $\ell_{\infty}$ norm and that under the $\ell_{2}$ norm, Eisenbrand [4] cites Lagarias’ work [5] to demonstrate that reduced lattice basis containing the $\ell_{\infty}$-shortest vector can be reduced to the $\ell_{2}$-shortest vector through a constant number of Lagrange reductions.
@@ I-B Motivation and Our Contribution
Existing research indicates that the SVP (SBP) in two-dimensional lattices remains imperfectly resolved. Under the commonly used Euclidean metric, Although Yap’s Half-Gaussian algorithm [15] achieves lower complexity than Lagrange reduction, its underlying fundamental reduction algorithm (CRS) – as Yap himself acknowledges – is not inherently more efficient than classical Lagrange reduction. This naturally raises a pivotal question: Does there exist a fundamental reduction algorithm that is more efficient than the Lagrange reduction algorithm, along with an optimized version that can outperform the Half-Gaussian algorithm in finding the $\ell_{2}$-shortest vector (basis)? For the special Minkowski metric, current algorithms face the following limitations:
  • For HNF-form bases $\mathbf{B}=\begin{pmatrix}a_{1}&b_{1}\\ 0&b_{2}\end{pmatrix}$: While prior works suggest that integer HGCD algorithms can optimize continued fraction-based methods, detailed technical analyses and implementations remain scarce. As noted in [11], “the integer HGCD algorithm is intricate, error-prone, and rarely fully detailed”, and [12] emphasizes its implementation challenges due to “numerous sub-cases and limited practical adoption.” Notably, [3] provides a concrete HGCD-based SVP solution for the special case $b_{2}=1$, but the implementation details for general $b_{2}$ is not available.
  • For general-form bases: Current algorithms first convert inputs to HNF via (HGCD-based) extended Euclidean algorithms. However, this conversion often inflates integer sizes, especially for short initial bases, leading to longer lattice vectors. This raises two interesting questions: (i) Is HNF conversion necessary? (ii) Can direct reduction algorithms bypass this step for a faster performance?
To address these challenges, this work aims to develop a novel fundamental reduction algorithm achieving: (1) For Euclidean metric: Superior efficiency over both Lagrange reduction and Yap’s CRS, with an optimized version surpassing Half-Gaussian. (2) For Minkowski metric: Elimination of HNF conversion and generalization of HGCD-based optimization to arbitrary $b_{2}$, thereby bridging the gap between theoretical possibilities and practical implementations. Concretely, our main contributions can be summarized as follows:
@@
    Novel Definition of a Reduced Basis. For the first time, we explicitly propose a novel definition of a reduced basis distinct from the conventional Lagrange-reduced basis. This newly defined basis not only contains the $\ell_{\infty}$-shortest vector but also yields the $\ell_{2}$-shortest vector in at most one additional addition/subtraction step. This framework extends the theoretical foundation of lattice reduction and is expected to provide a solid foundation for further research in this field.
@@
    Efficient Algorithms with Detailed Implementation and Rigorous Complexity Analysis. For our newly introduced reduced basis, we design a novel fundamental algorithm named CrossEuc that bypasses HNF conversion, achieving a complexity of $O(n^{2})$. This approach eliminates the computational overhead associated with HNF transformation. Furthermore, we introduce a vectorized adaptation of the HGCD algorithm, termed HVec, which represents the first explicit extension of HGCD-like techniques to vector pairs. This innovation is of independent theoretical interest and significantly advances the efficiency of solving the SVP in two-dimensional lattices. Building on this, we present an optimized version HVecSBP of the fundamental algorithm CrossEuc that reduces the complexity to $O(\log n\cdot M(n))$, where $M(n)$ denotes the cost of multiplying two $n$-bit integers.
@@
    Comprehensive Experimental Performance Analysis. We conduct extensive experiments comparing our methods against state-of-the-art algorithms. Results show that (1) For computing the $\ell_{2}$-shortest basis with input lattice basis size $2\times 10^{5}$, regardless of whether the input is in HNF form or general form, our proposed fundamental reduction algorithm, $\mathbf{CrossEuc}$, achieves approximately 500×500$\times$500 × and 350×350$\times$350 × speedup compared to Yap’s foundamental reduction algorithm $\mathbf{CRS}$ and the classical Lagrange reduction algorithm, respectively. In comparison with Yap’s optimized algorithm, $\mathbf{HalfGaussianSBP}$, our optimized variant, $\mathbf{HVecSBP}$, achieves an additional speedup of approximately 14×14$\times$14 ×. (2) For computing the $\ell_{\infty}$-shortest basis with input lattice basis size $10^{6}$, our new algorithms demonstrate significant improvements: (2.1) For HNF-form bases, $\mathbf{HVecSBP}$ achieves 13×13$\times$13 × acceleration over prior designs in worst-case scenarios. (2.2) For general-form bases, compared to the previous designs without using HGCD optimization, the proposed algorithm CrossEuc achieves a at least 4.5×4.5$\times$4.5 × efficiency improvement. Compared to the previous designs using HGCD optimization, when the input vectors are nearly degenerate (i.e., treated as integers), HGCD-HNF-HVecSBP method retain a slight advantage. As the linear dependency between vectors weakens, HVecSBP exhibits remarkably growing efficiency gains.
@@ II-A Notations and Terminologies
We use $M(n)$ to denote the time complexity of the multiplication of two $n$-bit integers. Uppercase bold letters represent matrices, while lowercase bold letters represent vectors. For any two vectors $\mathbf{a}=(a_{1},a_{2})^{T}$ and $\mathbf{b}=(b_{1},b_{2})^{T}$, we define #$\#\mathbf{a}=\max\{\#a_{1},\#a_{2}\}$, #$\#(\mathbf{a},\mathbf{b})=\max\{\#\mathbf{a},\#\mathbf{b}\}$ and #¯$\underline{\#}(\mathbf{a},\mathbf{b})=\min\{\#\mathbf{a},\#\mathbf{b}\}$. For a vector $\bm{x}=(x_{1},x_{2},\ldots,x_{m})\in\mathbb{R}^{m}$, its Euclidean length ($\ell_{2}$ norm) is defined as $\|\bm{x}\|=\sqrt{\sum_{i=1}^{m}x_{i}^{2}}$, and its Minkowski length ($\ell_{\infty}$ norm) is defined as $\|\bm{x}\|_{\infty}=\max_{1\leq i\leq m}|x_{i}|$.
@@ III Our new reduced basis and its properties
In this section, we first introduce a new reduced basis and then present its properties. Throughout this section and the rest of the paper, unless otherwise specified, the notations $\|\mathbf{x}\|$ and $\|\mathbf{x}\|_{2}$ are used to represent the Minkowski metric ($\ell_{\infty}$ norm) of $\mathbf{x}$ and the Euclidean metric ($\ell_{2}$ norm), respectively.
@@
Let $\mathcal{L}=L(\bm{B})$ be a lattice with basis $\mathbf{B}$ defined in equation (1), and let $\lambda_{1}(\mathcal{L})$ and $\lambda_{2}(\mathcal{L})$ denote the successive minims under $\ell_{2}$ norm. If $\mathbf{B}$ is reduced, then $\mathbf{u}=\mathrm{argmin}\{\|\mathbf{a}\|_{2},\|\mathbf{b}\|_{2},\|\mathbf{a}+\mathbf{b}\|_{2},\|\mathbf{a}-\mathbf{b}\|_{2}\}$ is a shortest non-zero lattice vector under the $\ell_{2}$ norm. That is, $\lambda_{1}(\mathcal{L})=\|\mathbf{u}\|_{2}$. Further, for any vector $\mathbf{x}\in\{\mathbf{a},\mathbf{b}\}\setminus\{\mathbf{u}\}$, $\mathbf{v}=\mathbf{x}-q\mathbf{u}$ with $q=\left\lceil\frac{\langle\mathbf{x},\mathbf{u}\rangle}{\langle\mathbf{u},\mathbf{u}\rangle}\right\rfloor$ is a lattice vector achieving $\lambda_{2}(\mathcal{L})$. That is, $\lambda_{2}(\mathcal{L})=\|\mathbf{v}\|_{2}$.